OpenStudy (anonymous):

The height of a rectangular wall is 8 meters less than its base. The base is 17 meters. What is the height of the wall? What is its area?

OpenStudy (anonymous):

H=17-8 A=(17-8)*(17)

OpenStudy (anonymous):

so height=11m, area=187m2